OpenStudy (anonymous):

The given measurements may or may not determine a triangle. If not, then state that no triangle is formed. If a triangle is formed, then use the Law of Sines to solve the triangle, if it is possible, or state that the Law of Sines cannot be used. B = 126°, c = 7, b = 12

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\frac{\sin(C)}{c}=\frac{\sin(B)}{b}\]